July 15, 2015 Calibration
GlueX Calibration Meeting
Wednesday, July 15, 2015
11:00 am, EDT
JLab: CEBAF Center, F326
You can connect using BlueJeans using the meeting number 630 804 895 . (Click "Expand" to the right for more details -->):
- Make sure you have created a BlueJeans account via your JLab CUE account using this link:
- http://jlab.bluejeans.com (You should only need to do this once)
- Meeting ID: 630804895
- (you may need to type this in, depending how you connect)
- If connecting via Web Browser: click this link (no passcode is needed):
- If connecting via iOS or Android App:
- Use your JLab e-mail address to log in and then enter the meeting ID given above to join the meeting
- If connecting via Phone: Dial one of the following numbers and then enter the meeting ID above and hit "#" or "##"
- Dial: 1 888 240 2560 (US or Canada only)
- or, List of International Numbers
- If connecting via Polycom unit:
- Dial 22.214.171.124 or bjn.vc
- Enter meeting ID above
- Use *4 to unmute
Talks can be deposited in the directory
/group/halld/www/halldweb/html/talks/2015 on the JLab CUE. This directory is accessible from the web at https://halldweb.jlab.org/talks/2015/ .
- Calibration Processing
- Potential uses of git?
- Run quality criteria
- Summary of Detector Performance
- Subdetector Reports
- Action Item Review
- Produce updated calibration plots and performance numbers, preferably within a week.
- Polish calibration routines, including testing of results, documentation (GlueX note).
- Update simulation parameters.
Attending: Sean (NU); Simon, Beni, Mark I., Justin, Eric, Lubomir, Paul, Eugene, Adesh, Mark D. (JLab); Curtis, Will M., Mike S. (CMU)
- S&T Review
- Plots from all of the detectors have been collected and put on both DocDB and the plot database in the GlueX portal.
- Justin is putting together his talk, and will send out requests for more information as needed.
- Latest monitoring run
- The most recent offline monitoring run was started last Friday. Database problems delayed processing of the results, but they should be done today. REST files are available as described in the email linked above.
- moving to git
- The repositories for sim-recon and HDDS are moving from Subversion to git. Basic instructions on how to interact with the new system are linked to above. If you plan to contribute code, you will need an account on github; if you just plan on using the latest code for analysis, you can just check out the main branch as described in the instructions.
- Note that before the move, all sim-recon and online plugins have been moved into a new directory in sim-recon: src/plugins
- A new directory for storing DANA calibration plugins has been created in sim-recon: src/plugins/Calibration
Sean has been testing this system in the past few weeks, and plans to make the first "official" run next week. He is shooting for Tuesday, and will make any annoucements related to this on the offline software mailing list.
Since a lot of progress has been made in calibration procedures, Sean went through a few guidelines that should be followed to standardize them:
- DANA plugins should be stored in the sim-recon directory src/plugins/Calibration
- Other scripts should be stored in some common area
- Suggestion: One directory per subdetector, organization in each directory is up to the subdetector groups
- Version control: Either a common location in SVN or individual git repositories. Feedback from detector groups would be helpful.
- The calibration jobs execute ROOT scripts in a common environment where libraries (compiled or ROOT scripts) can be pre-loaded.
- Quality testing routines should be included for each calibration procedure
- All output should be to files, not the screen, so that results can be properly staged to disk
- Documentation: both how to run scripts and GlueX note draft.
Sean also suggested that detector groups start coming up with data quality criteria so that we can start working on automatically classifying run quality. There was a lot of discussion on what this would mean, and Sean offered to come up with a more concrete plan that could be further discussed. Some points that were brought up:
- Subdetectors can have transitory problems. We might want to track these (in the CCDB perhaps?), and then map these to lists of files which should be excluded.
- In the current software, calibrations can only be done on a per-run basis. Changing to a finer grained system would involve a lot of work.
- Determination of pedestals on a per-file basis is desirable, based on the data collected so far.
- Mark D. shared his experience from QWeak. They took long runs and did calibrations/data quality on a per-file basis, since they found they were losing ~6% of run time due to starting and stopping CODA. Eugene pointed out that the newer CODA we are using is much faster than the old one.
Summary of Detector Performance
Curtis has started collecting information on the design and current values of detector performance parameters (e.g. resolutions and efficiencies). The intent is to have a snapshot of detector performance.
Detector groups are requested to send Curtis updated values. The status of the document will be reviewed at the next meeting.
Adesh discovered that the non-linear corrections he added with the updated gains last Friday did not entirely fix the pi0 peak location to the expected location, and backed out these corrections before the monitoring run started. He is investigating it further.
Elton looked at determining gains with electrons matched to the FCAL using data taken with the magnetic field on, and found the gain factors to be ~20% higher (i.e. lower gains) than those obtained from data taken with the magnetic field off. The pi0 peak is correspondingly lower with the magnetic field on. This effect needs to be both properly calibrated and understood.
No news - most of the current effort is going into determining specifications for new fADC250 firmware.
Mike S. showed us some preliminary slides he is putting together. There were two main topics:
- Alignment using Millipede
- After studying a regularization parameter, he has been able to determine that his procedure works for simulated data with gaussianly distributed wire offsets with a sigma of 150um.
- If the sigma of the offset distribution is increased to 300um, then the procedure does not work as well, and has a tendency to pick up on statistical fluctuations.
- By changing the regularization parameter and iterating the procedure, he is able to get better results for this case, and is working on further improvements
- Potential straw deformation
- It has been noted since last year that the drift time distributions for many straws do not have a clean endpoint as expected.
- Mike has been investigating obtaining the drift time/distance relation from data, instead of using results from Garfield calculations, and in the long cosmic runs from the spring noticed that many straws seemed to have two different endpoints.
- Looking at the drift time distributions as a function of the azimuthal angle of the track in the straw shows a dependence, suggesting that the straws are deformed.
- This effect needs to be understood in detail, but for now a correction in local z and phi coordinates can be done. A much larger sample of cosmics is needed. Currently Mike is using data from a large run over a weekend, roughly 24-36 hours according to Beni. Roughly an order of magnitude more data is needed.
This will be discussed in more detail in the drift chamber meeting tomorrow.
Lubomir is working on implementing corrections to the centroid calculation using the cathode strips. There is a roughly sinusoidal correction, relative to the wire positions. In preliminary tests, he finds that these corrections are about a factor 4 too small to correct for the larger than expected resolutions. Another possible cause could be the size of the avalanches, which depends on the gas mixture.
He and Luke are taking data using the iron source with the test chamber to understand these effects.
Beni's latest calibrations have been added to the CCDB, and he has started a technical note to describe them, which can be found here. His calibrations have only been done with TDC data - the next challenge is to calibrate the ADC's.
Sean pointed out that initial monitoring results show a greatly improved efficiency for tracks being matched in the TOF.
The time resolution with respect to the RF has been measured to be 300 ps, which is 50 ps better than the design goal! The effect of the propogation time corrections in reconstruction has been studied, and the results are found to have little difference than the hard-coded value currently used.
Eric is working on rechecking the time resolution for tracks with a match in the TOF.
Alex B. showed some preliminary results for his attempt to determine corrections for the PS energy scale. A significant improvement is seen, and he is continuing to refine these results.
Nathan is back from vacation, and digging back into the data.
Mark, Justin and Sean have been discussing ways to better monitor the results of the simulations. They came up with a list of plugins to be run over the simulations. They also decided to not create or analyze EVIO files. Mark will restart the simulations with these conditions, hopefully by the end of the day.